You are in charge of erecting a radio telescope on a newly discovered planet. To minimize interference, you want to place it where the magnetic field of the planet is the weakest. The planet is spherical, with a radius of 6 units. Based on a coordinate system whose origin is the center of the planet, the strength of the magnetic field is given by

M(x,y,z)= 09 + 22 + z - 19

Where should you locate the radio telescope?
Answer #1

solution: of a Given that the equation of the surface sphere of radius is 6 Magnetic field M(x, y, z)= 6x-4² + 2z+60 Let Centthe second equaton tells that have و طرح حنه می ره تو (اپ 2) - Y=0 on a = -1 IF 2--/ then the other equators are w: -2Z & 6+6+2=40²2 6 = 4a?z-z 6 = (42²-,)z Z : - n-- 72 n= 122 42², Note that we can divide by 42²-1 bells us that it cannot equal 2000- 12 03 + 2 (» در به) - ( بهای کے 12 :(2 را به سر من 603+ (= - - - - ( مداد کی 29 = (2 ,4 . --| | 603 + ریه - رده 60 = (0 ,0